40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 28 02:43:49 2013, yuta, Update, LSC, some qualitative evidence of PRMI sideband lock PRMIsideband.png
    Reply  Thu Feb 28 03:25:27 2013, yuta, Update, LSC, PR gain ~ 25 from PRMI carrier lock PRMIcarrier.png
       Reply  Thu Feb 28 03:41:31 2013, Koji, Update, LSC, PR gain ~ 25 from PRMI carrier lock 
       Reply  Sat Mar 2 05:53:15 2013, yuta, Update, LSC, stable lock of PRMI LSCMICHOLTF_PRMI.pngLSCPRCLOLTF_PRMI.pngPRMtoAS55REFL55.png
          Reply  Sat Mar 2 14:52:02 2013, rana, Update, LSC, stable lock of PRMI 
             Reply  Sun Mar 3 22:16:46 2013, Jenne, Update, LSC, LSC whitening triggering started 
                Reply  Mon Mar 4 09:55:33 2013, rana, Update, LSC, LSC whitening triggering started 
                Reply  Tue Mar 5 18:36:27 2013, Jenne, Update, LSC, LSC whitening triggering started 
Message ID: 8217     Entry time: Mon Mar 4 09:55:33 2013     In reply to: 8215
Author: rana 
Type: Update 
Category: LSC 
Subject: LSC whitening triggering started 

  How about posting a logic flow diagram? Is the idea to trigger only on the power signals to determine the lock state? Is the hysteresis going to be done in the same way as the main filter bank triggers?

ELOG V3.1.3-